# Predictability-Aware Compression and Decompression Framework for Multichannel Time Series Data

### v1

@@ Abstract
Real-world multichannel time series prediction faces growing demands for efficiency across edge and cloud environments, making channel compression a timely and essential problem. Motivated by success of Multiple-Input Multiple-Output (MIMO) methods, we propose a predictability-aware compression–decompression framework to reduce runtime, lower communication cost, and maintain prediction accuracy across diverse predictors. The core idea involves using a circular periodicity key matrix with orthogonality to capture underlying time series predictability during compression and to mitigate reconstruction errors during decompression by relaxing oversimplified data assumptions. Theoretical and empirical analyses show that the proposed framework is both time-efficient and scalable under a large number of channels. Extensive experiments on six datasets across various predictors demonstrate that the proposed method achieves superior overall performance by jointly considering prediction accuracy and runtime, while maintaining strong compatibility with diverse predictors.
1 Introduction
Multichannel time series prediction [21, 32, 6, 10, 2, 22] is widely studied across various application domains, including but not limited to manufacturing process monitoring [11], traffic flow analysis [13], energy consumption tracking [29], and environmental sensing [20], where a channel usually stands for a dimension (e.g., latitude in location data) or a modality (e.g., temperature in environmental sensing). These applications typically stream data from edge-device sensors to cloud-based prediction services, posing practical challenges for transmission and computation. As shown in Figure 1, real-world deployments often involve long time series data with dozens or even hundreds of channels, which strain communication bandwidth and computational resources on edge devices. While compressing along the temporal dimension may reduce input size, it can obscure crucial short-term patterns and make downsampling a potentially risky operation for prediction tasks. These situations make channel compression for time series prediction a timely and valuable research direction.
Figure 1: Application requirements in the real world multichannel prediction
@@ 1 Introduction
Despite the importance of channel compression, research on multichannel compression for time series prediction remains limited. General-purpose dimension reduction techniques, including Principal Component Analysis (PCA) [9], Multi-Dimensional Scaling (MDS) [27], Isomap [26], t-distributed Stochastic Neighbor Embedding (t-SNE) [15], and Uniform Manifold Approximation and Projection (UMAP) [16], remain commonly applied in time series prediction tasks. For example, DeepGLO [25] employs PCA-based matrix factorization to extract global features, while another paper [8] applies PCA for temporal dimension reduction. Although these methods aim to capture either temporal inter-channel dependencies or perform intra-channel reduction, they are not specifically designed for time series prediction. As a result, they fail to integrate both inter- and intra-channel information in a unified manner, and do not preserve or reveal the underlying predictability of the data after compression. Additionally, many of these methods, originally developed for data visualization or non-regression tasks, are unable to accurately reconstruct the original multichannel structure from the predicted compressed representation.
@@
The Multiple-Input Multiple-Output (MIMO) paradigm in signal processing, along with the recent success of MIMONet [17] in image compression, motivates the design of a compression–decompression framework for multichannel time series prediction. During the design process, we identify the following challenges. First, preserving the potential predictive information is non-trivial in the compression process. The temporal dependence inherent in time series data stores predictive information and provides a foundation for accurate forecasting. However, such predictability cannot be guaranteed during compression when using current approaches, such as circular convolution with keys sampled from a normal distribution (as in HRR [23]) or a Bernoulli distribution (as in MIMONet). Second, oversimplified assumption poses a critical challenge for accurate decompression. Specifically, the assumption of a normal distribution, as adopted in HRR, does not align well with the characteristics of time series data. This mismatch leads to inappropriate decompression module and additional errors, including intra-channel interference across timestamps and inter-channel interference within the same timestamp. Third, it is essential to jointly consider the compression-decompression and prediction processes, and explore better overall end-to-end performance. While compressed data naturally leads to reduced computational and transmission time, the compression-decompression steps themselves may introduce additional overhead and information loss, potentially increasing overall time complexity and degrading prediction performance.
To address these challenges, we propose a predictability-aware compression–decompression framework for multichannel time series prediction. In the compression stage, periodicity keys, derived from the least common period across channels via FFT or domain knowledge, are used in a circular convolution process to preserve the predictability information. To mitigate reconstruction errors, the decompression stage employs a circular Toeplitz key matrixs and a residual block. To ensure time efficiency, we provide theoretical analysis and introduce sparse encoding and lightweight modules to reduce overhead, particularly under long sequences or edge-device constraints. Extensive experiments on six datasets and multiple backbone predictors (transformer, linear, MLP, and convolutional) demonstrate that our framework consistently improves runtime efficiency and keep prediction accuracy. We also introduce a Cobb–Douglas-based [4] performance index to jointly evaluate accuracy and efficiency, showing that our method achieves superior overall performance and strong compatibility across predictors.
Based on the above descriptions, our main contributions are summarized as follows:
  1. 1.
    A predictability-aware compression–decompression framework is proposed for multichannel time series prediction, which preserves data predictability during compression and minimizes reconstruction error during decompression.
  2. 2.
    Theoretical analyses demonstrate the time efficiency and scalability of the proposed method under a large number of channels, further supported by sparse encoding and lightweight residual components.
  3. 3.
@@
    Extensive experiments demonstrate that the proposed framework achieves lower runtime and superior performance when jointly considering prediction accuracy and computational cost, while maintaining strong compatibility across diverse module architectures.
2 Methodology
Figure 2: Compression and decompression framework with computational complexity annotations. The predictor (center) is not part of our proposed design and can be replaced by standard single-channel predictors.
@@ 2 Methodology
As shown in Figure 2, the overall framework consists of three components: a compression module, a decompression module, and a predictor. In the compression module, a multichannel time series input $X_{t-L:t}^{C}\in\mathbb{R}^{L\times C}$ is transformed into a compressed single-channel representation $X_{t-L:t}\in\mathbb{R}^{L\times 1}$, where $L$ denotes the historical length and $C$ is the number of channels. This single-channel representation is then passed through the predictor to generate the next $H$ time steps, denoted as $X_{t:t+H}\in\mathbb{R}^{H\times 1}$. The predicted sequence is subsequently decompressed to reconstruct the multichannel output $X_{t:t+H}^{C}\in\mathbb{R}^{H\times C}$.
This work primarily focuses on the compression and decompression modules. Specifically, we investigate how to effectively encode and retain relevant information during compression, and how to reconstruct the multichannel output more accurately by avoiding the oversimplified assumptions that often cause significant reconstruction errors in conventional methods. In the following sections, we sequentially detail the compression process, decompression process, computational complexity analysis, and the overall workflow during training and inference.
2.1 Compression process
The key distinction between the proposed multichannel compression framework and conventional compression methods lies in its predictability-awareness, which refers to the capability to preserve and uncover latent temporal dependencies within time series data. By maintaining these dependencies throughout the compression process, the framework remains compatible with a wide range of predictors, including MLP, linear, convolutional, and transformer-based models that rely on temporal continuity and structure.
In this section, we first define a time series predictability index, followed by the introduction of a predictability-aware compression mechanism, which includes a single-channel encoding strategy and a multichannel combination process.
Time series data predictability
Correlation serves as a central metric for identifying predictive patterns in temporal data. For example, Autoformer [31] replaces the attention mechanism with an auto-correlation module to model temporal dependencies, while Corad [12] leverages Pearson correlation to compress similar segments.
In this work, we define predictability as the degree to which future values depend on historical observations. Following prior research, we adopt the Pearson correlation coefficient to quantify the predictability in time series data. The coefficient, ranging from 0 to 1, measures the strength of the linear association between the former sequence and the current sequence.
Single channel enhancement
Motivated by Holographic Reduced Representations (HRR), a circular convolution process is employed to encode multichannel time series data. To preserve and reveal the potential predictability of each channel, seasonal keys are proposed to guide the compression process. The seasonal key construction and the channel encoding formulation are presented below:
$Y_{i}$ $=\text{Encode}(K^{i},X^{i})$ (1)
$=[\sum_{z=0}^{n-1}x_{z}^{i}\,k_{(0-z)},\sum_{z=0}^{n-1}x_{z}^{i}\,k_{(1-z)},...,\sum_{z=0}^{n-1}x_{z}^{i}\,k_{(l-z)}]$
$s.t.\quad k_{t}=k_{t+\tau}$

### v2

@@ Abstract.
Real-world multichannel time series prediction faces growing demands for efficiency across edge and cloud environments, making channel compression a timely and essential problem. Motivated by success of Multiple-Input Multiple-Output (MIMO) methods in signal processing, we propose a predictability-aware compression–decompression framework to reduce runtime, decrease communication cost, and maintain prediction accuracy across diverse predictors. The core idea involves using a circular seasonal key matrix with orthogonality to capture underlying time series predictability during compression and to mitigate reconstruction errors during decompression by introducing more realistic data assumptions. Theoretical analyses show that the proposed framework is both time-efficient and accuracy-preserving under a large number of channels. Extensive experiments on six datasets across various predictors demonstrate that the proposed method achieves superior overall performance by jointly considering prediction accuracy and runtime, while maintaining strong compatibility with diverse predictors.
Data compression, Deep learning, Time series prediction
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792488isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Neural networksccs: Computer systems organization Sensor networksccs: Information systems Web services
1. Introduction
Multichannel time series prediction is widely studied across various domains (Jang and others, 2025; Liu and others, 2023; Tu and others, 2024; Nguyen and others, 2024) and has become increasingly critical in Web-scale applications, where time series data is transmitted through Internet of Things (IoT) networks to Web platforms for large-scale access and decision support. In this context, a channel refers to a single signal stream or time-dependent variable—such as the reading from one sensor, a specific feature dimension, or a data source—recorded in parallel with other channels as part of a multichannel time series. These applications typically stream data from sensors to Web platforms, posing practical challenges for transmission and computation. For example, in large-scale IoT deployments, low-cost sensors rely on edge devices to aggregate local data and transfer it from Low Power Wide Area (LPWA) networks to IP link network (Milarokostas et al., 2023). In this case, as shown in Figure 1, multichannel data may originate from a single sensor (Raza et al., 2017) or from edge devices aggregating multiple sensors in areas without IP coverage (Ghosh et al., 2023; Wu et al., 2023), both of which strain bandwidth and computation. Therefore, time series data compression for prediction task is a timely and valuable research direction.
Figure 1. Time series data compression in two example situations.
None.
Existing time series data compression methods primarily focus on storage and computational efficiency through temporal reduction, which can be categorized into three types. (1) Representation-based methods (Li et al., 2014; Chandak et al., 2020; Moon et al., 2022) use mathematical functions or learned nonlinear mappings to obtain compact latent representations. (2) Dictionary-based methods (Lillo-Castellano et al., 2013; Ryabko, 2009) exploit structural redundancy via similarity measures or codebook matching. (3) Sequential algorithms (Ryabko et al., 2016; Lysyak and Ryabko, 2016; Lillo-Castellano et al., 2013; Ryabko, 2009) process data incrementally using information-theoretic principles for online or model-free prediction. While compressing along the temporal dimension reduce input size, it can obscure crucial short-term patterns and make down-sampling a potentially risky operation for prediction task (Zhu et al., 2025).
@@ 1. Introduction
On the other hand, general-purpose dimension-reduction techniques can be directly applied for channel compression, including Principal Component Analysis (PCA) (Hotelling and others, 1933), Multi-Dimensional Scaling (MDS) (Torgerson and others, 1952), Isomap (Tenenbaum and others, 2000), t-distributed Stochastic Neighbor Embedding (t-SNE) (van der Maaten and Hinton, 2008), and Uniform Manifold Approximation and Projection (UMAP) (McInnes and others, 2018). However, while these methods aim to either capture temporal inter-channel dependencies or perform intra-channel reduction, they are not specifically designed for time series prediction. As a result, they fail to integrate both inter- and intra-channel information in a unified manner, and do not preserve or reveal the underlying predictability of the data after compression. Additionally, many of these methods, originally developed for data visualization or non-regression tasks, are unable to accurately reconstruct the original multichannel structure from the predicted compressed representation.
@@
The Multiple-Input Multiple-Output (MIMO) paradigm in signal processing and the recent success of MIMONet (Menet and others, 2023) in image compression bring new opportunities to build a compression-decompression framework for multichannel time series prediction. However, applying this idea in practice still poses significant challenges. First, preserving the potential predictive information is non-trivial in the compression process. The temporal dependence inherent in time series data stores predictive information and provides a foundation for accurate forecasting. However, such predictability cannot be guaranteed during compression when using current approaches, such as circular convolution with keys sampled from a normal distribution (as in HRR (Plate, 1995)) or a Bernoulli distribution (as in MIMONet). Second, oversimplified assumptions on data pose a critical challenge for accurate decompression. In particular, the assumption of normally distributed input—adopted in HRR—does not align well with the properties of real-world time series data. This mismatch can result in inadequate decompression modules and additional errors, such as intra-channel interference across timestamps and inter-channel interference within the same timestamp. Third, the overhead introduced by the compression–decompression processes may offset its efficiency benefits. While compressed data naturally leads to reduced computational and transmission time, the compression-decompression steps themselves may introduce additional overhead, potentially increasing overall time complexity.
To address these challenges, we first theoretically show that redundancy in multichannel time series can be exploited to design a compression method that preserves predictive information. Based on this insight, we propose a predictability-aware compression-decompression framework for multichannel time series prediction. The predictability is maintained by extracting the underlying seasonal patterns through Fourier transform. In this framework, the compression module uses a circular-key for single channel encoding and multichannel predictability alignment, while the decompression module applies the same circular-key matrix with orthogonality constraints and a residual block to mitigate reconstruction errors. Time-complexity analysis confirms its superior efficiency over direct multichannel predictors. Extensive experiments on six datasets and four backbone predictors demonstrate runtime reduction without sacrificing accuracy, even for data without strong periodicity (e.g., sensor drift data). A Cobb–Douglas-based performance index (Cobb and others, 1928) jointly evaluates accuracy and efficiency, highlighting our framework’s overall advantage and broad compatibility across diverse predictors.
Our main contributions are summarized as follows:
  • Observations and Theoretical Foundations: We provide theoretical analyses that prove the predictability-awareness of our compression module, grounding the results in empirical observations and supporting analysis.
  • Compression–Decompression Framework: We propose a compression–decompression framework for multichannel time-series prediction that minimizes reconstruction error during decompression. Time-complexity analysis shows that the framework achieves better time efficiency compared with a direct multichannel predictor.
@@
  • Experimental Analysis: We conduct extensive experiments demonstrating that the proposed framework achieves lower runtime and superior performance when jointly considering prediction accuracy and computational cost, while maintaining strong compatibility across diverse module architectures.
The following sections are organized into Related Work, Problem Definition, Observations and Theoretical Foundations, Compression and Decompression Framework, and Experiments. The Observations and Theoretical Foundations section shows how our method preserves predictability. The Compression and Decompression Framework section explains the details shown in Fig. 3 and describes overall process in detail.
2. Related Works
Time Series Prediction
Deep learning architectures for multichannel time series forecasting can be broadly categorized into several paradigms. Channel-independent models (Nie and others, 2023; Zeng and others, 2023) treat each channel separately, ignoring inter-series correlations to achieve scalability and simplicity. In contrast, channel-mixing models (Ekambaram and others, 2023; Huang and others, 2024; Cai and others, 2024) explicitly model dependencies across different channel to improve predictive accuracy. Single-channel architectures (Palaskar and others, 2024) are originally designed for single channel inputs but can be generalized to multichannel settings. In addition, recent efforts have explored using large pre-trained models for time series forecasting. Time-LLM (Jin et al., 2024) reprograms frozen large language model (LLM) by encoding temporal data as token sequences to enable zero-shot forecasting. UniST (Yuan et al., 2024) proposes a universal prompt-based spatio-temporal predictor that adapts to a wide range of urban forecasting tasks.
The growing use of deep learning and LLM in time series forecasting, which requires massive volumes of data, calls for compression methods that preserve predictive information under edge-device and sensor constraints.
Time Series Data Compression
There are two main aspects of time series compression: one along the temporal dimension and the other along the channel dimension. Existing time-series compression methods primarily focus on storage and computational efficiency through temporal reduction, which can be categorized into three types. (1) Representation-based methods (Li et al., 2014; Chandak et al., 2020; Moon et al., 2022) use mathematical functions or learned nonlinear mappings to obtain compact latent representations. (2) Dictionary-based methods (Lillo-Castellano et al., 2013; Ryabko, 2009) exploit structural redundancy via similarity measures or codebook matching. (3) Sequential algorithms (Ryabko et al., 2016; Lysyak and Ryabko, 2016; Lillo-Castellano et al., 2013; Ryabko, 2009) process data incrementally using information-theoretic principles for online or model-free prediction. On the other hand, general-purpose dimension-reduction techniques—such as Principal Component Analysis (PCA) (Hotelling and others, 1933), Multi-Dimensional Scaling (MDS) (Torgerson and others, 1952), Isomap (Tenenbaum and others, 2000), t-distributed Stochastic Neighbor Embedding (t-SNE) (van der Maaten and Hinton, 2008), and Uniform Manifold Approximation and Projection (UMAP) (McInnes and others, 2018)—are often applied for channel compression without task-specific design. For example, DeepGLO (Sen and others, 2019) adopt low-rank matrix factorization inspired by the principles of PCA to extract global features, although not specifically for time series data compression.
Multiple-Input Multiple-Output (MIMO)
Techniques such as MIMO with Sparse Code Multiple Access (Han et al., 2016; Kosasih et al., 2018), Non-Orthogonal Multiple Access (Ding and Cai, 2017; Wang et al., 2020), and Code Division Multiple Access (Liao, 2007; Boukenadil et al., 2024) have been proposed to allow different signals to share the same communication channel. A key principle in their design is to ensure that signals remain distinguishable and do not interfere with each other during transmission, typically through coding, modulation, or power allocation strategies. However, time series prediction requires preserving temporal context and inter-channel complementary information under compression transmission, which these methods do not support. Thus, while we are inspired by their channel sharing strategies, they are not directly applicable to predictive tasks.
3. Problem Definition
Before presenting the design of our method, we first describe the format and assumptions of the input multichannel time series data. We then introduce the objective function that guides our compression and prediction framework.
Definition 1:
The number of input channels $C\in\mathbb{N}$ of the input time series data $X\in\mathbb{R}^{L\times C}$ refers to the input dimension that is distinct from the temporal axis with number of timestamps $L$.
Assumption 1:
The multi-channel time series input $X_{t-L:t}^{C}\in\mathbb{R}^{L\times C}$ contains latent seasonal patterns, and the input length $L$ is assumed to be longer than the most dominant seasonal period. Here, “seasonal period” denotes the dominant frequency component (e.g., via FFT), not a strict requirement for an obvious seasonal pattern in the data.
Objective Function
Motivated by the Cobb–Douglas production function (Cobb and others, 1928), we define the objective function as a multiplicative function of prediction accuracy and runtime, as shown below:
(1) $\min_{\theta}$ $\mathcal{F}_{\text{pred}}\left(\hat{X}_{t:t+H}^{C},\ X_{t:t+H}^{C}\right)\cdot\mathcal{F}_{\text{time}}\left(\theta\right)$
s.t. $\operatorname{dim}\left(\text{Compressed}(X_{t-L:t}^{C})\right)=(L,1)$
In this formulation, $\theta$ denotes a set of parameters corresponding to the compression, prediction, and decompression modules. The variables $\hat{X}_{t:t+H}^{C}$ and $X_{t:t+H}^{C}$ represent the predicted and ground truth sequences for the next $H$ time steps. The term $\mathcal{F}_{\text{pred}}$ quantifies the prediction error (e.g., Mean Squared Error), while $\mathcal{F}_{\text{time}}$ represents the total inference latency. The constraint enforces that the compressed representation maintains a single channel, i.e., $\operatorname{dim}=(L,1)$.
@@ 3. Problem Definition
In this design as shown in Fig. 3, a multichannel time series input $X_{t-L:t}^{C}\in\mathbb{R}^{L\times C}$ is first compressed into a single-channel representation $X_{t-L:t}\in\mathbb{R}^{L\times 1}$. This compressed sequence is used to predict the next $H$ time steps, resulting in $X_{t:t+H}\in\mathbb{R}^{H\times 1}$, which is then decompressed to reconstruct the multichannel output $X_{t:t+H}^{C}\in\mathbb{R}^{H\times C}$.
4. Observations and Theoretical Foundations
In this section, we first identify the redundancy phenomenon in multichannel time series data. Based on this observation, we present Theorem 1, which reveals the opportunity to remove redundant information while retaining predictive information during the compression process. Building on this theoretical opportunity, we introduce Theorem 2 to support that our proposed method is predictability-aware.
4.1. Redundancy in Multichannel Time Series
Multi-channel time series data can be represented by a limited number of channels.
As shown in Table 1, the multichannel time series data across all datasets exhibit substantial redundancy. For example, in the Sensor Drift dataset, only 2 out of 321 principal components are sufficient to retain 95% of the total variance, and the top 50 components capture 100% of it. Similarly, NYC taxi and DC bike require fewer than 15 components to preserve 95% of the variance, with over 99% already explained by the top 50. Additionally, the first two components alone explain a dominant portion of the variance (e.g., 92.4% and 5.7% in Sensor Drift), indicating that the data lie near a low-dimensional manifold. These observations validate the presence of strong redundancy in the input channels and motivate the use of compact representations through channel compression.
Table 1. PCA summary across datasets. PCs @95% var.: Number of principal components needed to retain 95% of variance; Top 50 var.: Cumulative variance explained by top 50 PCs; PC1/PC2 var.: Variance explained by the 1st and 2nd principal components.
Metric NYC_taxi DC_bike Sensor Drift
PCs @95% var. 15 / 321 13 / 321 2 / 321
Top 50 var. 0.9926 0.9908 1.0000
PC1 var. 0.5422 0.5723 0.9242
PC2 var. 0.1525 0.1993 0.0569
(a) NYC_taxi
(b) DC_bike
(c) Sensor Drift
Figure 2. Heatmaps of raw multichannel time series data. The vertical axis represents 321 channels, and the horizontal axis denotes 321 time steps. Highlighted regions indicate examples of similar patterns.
Figure 3. Overall architecture of the compression–decompression framework. Compression and decompression processes take place at the sensor or edge, while prediction is performed in the cloud. The framework is compatible with different deep learning prediction modules, such as transformer-based, MLP-based, and linear-based modules.
Multi-channel time series data exists redundancy about similar pattern and low variation components.
As shown in Figure 2, the heatmaps of the raw multichannel time series data reveal clear structural patterns across different datasets. A substantial number of channels exhibit little to no variation over time, as evidenced by large uniform-colored regions. In addition, multiple channels demonstrate similar temporal evolution patterns (such as the red box in Figure 2), suggesting redundancy and high correlation across channels. For example, in the DC bike dataset, certain areas show very few rentals. However, these data cannot be simply discarded, since there are still occasional rentals at specific times, although rentals are absent for most of the time. In addition, similar patterns can be observed in the DC dataset, which may arise from the same functional zones of the city or from nearby stations. These observations reveal the existence of compressible structure in data and justify an effective channel compression strategy prior to downstream prediction. To enable tractable modeling of such redundancy, the following assumption is made:
4.2. Retaining Predictive Information
Assumption 2: Gaussian redundancy assumption in multichannel time series.
To model the redundancy or low-variation components often observed in multichannel time series data, we assume that these components follow a zero-mean multivariate Gaussian distribution. Specifically, the redundancy-induced uncertainty across all channels and time steps is represented as:
(2) $\varepsilon\sim\mathcal{N}(0,\ \sigma^{2}I_{CL}),$
where $\varepsilon\in\mathbb{R}^{CL}$ is the vectorized noise across all channels and time steps, $C$ is the number of channels, $L$ is the sequence length, and $I_{CL}$ is the identity matrix of size $CL\times CL$. Each per-channel noise vector $\varepsilon^{(i)}\in\mathbb{R}^{L}$ is assumed to be independent and identically distributed.
Motivated by Information Bottleneck (IB) (Hu et al., 2024), we have the following theorem on multi-channel time series.
Theorem 1: Upper bound on predictive information reduction under compression
Under Assumption 2, consider a compressed representation normalized to have unit signal variance. The IB admits the following upper bound:
(3) $\mathcal{L}_{\text{IB}}\leq\frac{1}{2}\log\left(1+\frac{1}{CL\cdot\sigma^{2}}\right).$
A detailed derivation is provided in the Appendix A.1.
Remark 1
This result highlights a fundamental trade-off: when the original number of channels $C$ or the sequence length $L$ is small, the upper bound on the IB loss becomes larger, indicating a higher risk of losing predictive information under compression. Conversely, larger $C$ or $L$ lead to a tighter bound, implying that high-dimensional inputs offer greater potential for safe and effective compression without sacrificing prediction accuracy. This insight theoretically supports the feasibility of compressing multichannel time series in scenarios with abundant channel or temporal resolution.
4.3. Predictability Awareness Compression
Circular convolution (Plate, 1995) provides a fixed-width vector operation for binding two representations and has been widely used as a method for compositional encoding in distributed representations. We treat this as a baseline mechanism for compression.
Definition 2: Correlation as a measure of predictability in temporal sequences.
In time series data, the Pearson correlation coefficient (ranging from 0 to 1) between a historical segment and a future segment reflects the degree of predictability (Wu and others, 2021; Khelifati and others, 2019), with higher correlation indicating stronger temporal dependency or clearly seasonal pattern. Specifically, for two aligned sequences $x,y\in\mathbb{R}^{L}$, a higher Pearson correlation coefficient implies that one sequence is more informative about the other.
Theorem 2: Circular keys preserve single channel predictability information.
Given Assumption 1 and Definition 2, when the circular key $K\in\mathbb{R}^{L}$ shares the same seasonal period as the latent dominant seasonal period of the input sequence $X_{t-L:t}^{C}$, the circular convolution can retain the seasonal alignment and thereby preserve predictive information in the compressed representation $Y$.
(4) $\begin{split}Y^{i}&=\text{Encode}(K^{i},X^{i})=X^{i}*K^{i},\\ &\text{s.t. }k_{t}=k_{t\bmod\tau},\quad\text{for some fixed }\tau\ll L,\end{split}$
where the key enforces a repeating structure with seasonal period $\tau$. The proof is provided in Appendix A.2.
Remark 2
A formal justification based on the Pearson correlation coefficient shows that when the seasonal key aligns with the latent seasonal period in the input, the encoded representation maximizes predictability, achieving a value of $R=1$.
Remark 3
From a signal processing perspective, the seasonal circular key acts as a matched filter, amplifying frequency components that correspond to the latent seasonality in the input. This alignment ensures that the compressed representation $Y$ emphasizes task-relevant information while filtering out redundant or irrelevant patterns.
Corollary 1: Inter-channel predictability alignment via Least Common Multiple Seasonal Period (LCMP)
Given Theorem 2, consider a multichannel input $\{X^{i}\}_{i=1}^{C}$, where each channel $X^{i}\in\mathbb{R}^{L}$ exhibits a latent dominant seasonal pattern with seasonal period $\tau_{i}$. To preserve inter-channel seasonal consistency and maintain the predictability of the combined representation, we define a shared seasonal period
(5) $\tau=\mathrm{lcm}(\tau_{1},\tau_{2},\dots,\tau_{C}),$
and constrain each circular key $K^{i}\in\mathbb{R}^{L}$ to satisfy $k^{i}_{t}=k^{i}_{t\bmod\tau}$. Each channel is then encoded as $Y^{i}=\text{Encode}(K^{i},X^{i})$, and the final compressed representation is the channel-wise summation:
(6) $Y=\sum_{i=0}^{C-1}Y^{i}=\sum_{i=0}^{C-1}\text{Encode}(K,X^{i}).$
Remark 4
This seasonal shared-period encoding ensures that all encoded sequences remain aligned, enabling effective cross-channel compression and preserving the overall temporal structure in the latent representation. The seasonal keys $K$ from different channels can be identical, since they follow the same seasonal period $\tau$. In the next part, we will introduce a sparse method with reduced compression time, with minor limitations in predictability.
5. Compression and Decompression Framework
Based on the theoretical analysis above, this section illustrates the process of compression and decompression, as well as the time complexity of sparse optimization. As shown in Fig. 3, the framework is simple and flexibly compatible with existing time series prediction modules. The compression process has been justified in the former section with two steps (this section introduces a more unified and sparse method), and the following decompression process also involves two error-removal steps. Then, we prove that the framework with a single-channel predictor has lower time complexity than a direct multichannel predictor. Additional details on the pseudocode and training strategy are provided in the Appendix A.4.
5.1. Compression
Based on the former explanation and justification about predictability protection compression, the compression process is shown below. A sparse compression strategy is introduced to further reduce the computational burden on edge devices.
Concretely, each channel $X^{i}\in\mathbb{R}^{L}$ is partitioned into non-overlapping fragments of length $\tau$, where $\tau$ is the least common period (LCP) across all channels. Only complete segments are encoded, and any remaining tail shorter than $\tau$ is omitted. The compressed output is computed as:
(7) $Y=\sum_{i=0}^{C-1}\sum_{j=0}^{M-1}\text{Encode}(X_{j\tau:(j+1)\tau}^{i},K),$
where $M=\left\lfloor\frac{L}{\tau}\right\rfloor$, and $K$ is the seasonal key shared across all segments. The time complexity of this sparse compression process is reduced from $\mathcal{O}(CL^{2})$ to $\mathcal{O}\left(\frac{CL^{2}}{\tau}\right)$. When the period is chosen as $\tau=\frac{L}{\log L}$, the complexity becomes $\mathcal{O}(CL\log L)$
5.2. Decompression
The decompression process is designed to mitigate the reconstruction errors introduced by the simplifying assumptions in HRR, particularly the assumption that input data follows a Gaussian distribution—an assumption that often does not hold for real-world time series.
Decoding process via circular correlation.
Inspired by HRR, the basic decoding process uses circular correlation with seasonal keys, defined as: